1.      Eight girls went to a pizza parlor. Each of them ate 1 ¾ slices of pizza. How many slices of pizza did they eat in all                                                                                

Operation/s:                                                                                                            

Answer:                                                                                                                       

Unit (Label)?

2.     Kyle used 1/8 cup of flour to make two hotcakes. How many cups of flour will he need to make a dozen of hotcakes?                                                          Operation/s:                                                                                                           Answer:                                                                                                                   Unit (Label)?3.     A student works for 3 3/5 hours a day in a printing shop. If he is scheduled to work on Monday – Friday, how many hours does he work in a week?Operation/s:                                                                                                          Answer:                                                                                                                   Unit (Label)?4.     Bill earned Php 450 in selling donuts. He spent 1/5 of the money on flowers for his mother’s birthday. How much of Bill’s money was left after buying the flowers?                                                                                                                  Operation/s:                                                                                                           Answer:                                                                                                                              Unit (Label)?5.     Mark climbed a palo sebo. The pole is 15 meters high. He already climbed 7/10 of the pole. How many meters did he need to climb to reach the top of the pole?            Operation/s:                                                                                                            Answer:                                                                                                                              Unit (Label)?  

1
Expert's answer
2021-11-09T09:47:48-0500

1.

Operation/s:   

"1\dfrac{3}{4}\cdot8=\dfrac{7}{4}\cdot8=14"


Answer: "14"

Unit (Label): slices of pizza.


2.

Operation/s:   

"\dfrac{1}{8}\cdot\dfrac{12}{2}=\dfrac{3}{4}"


Answer: "\dfrac{3}{4}"

Unit (Label): cups of flour.


3.

Operation/s:   

"3\dfrac{3}{5}\cdot5=\dfrac{18}{5}\cdot5=18"


Answer: "18"

Unit (Label): hours.


4.

Operation/s:   

"450-\dfrac{1}{5}\cdot450=360"


Answer: "360"

Unit (Label): Php.


5.

Operation/s:   

"15-\dfrac{7}{10}\cdot15=4\dfrac{1}{2}"


Answer: "4.5"

Unit (Label): meters.